40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 289     Entry time: Mon Nov 27 20:24:24 2017
Author: Koji 
Type: General 
Category: General 
Subject: A former LHO PD (Trans) removed from the OMC #002 for the shipment to Stockholm 

Attachment 1: The PD was removed from the transmission side of the OMC #002 (former LHO OMC - the one blasted by the optical pulse in Aug 2016).
It was confirmed that the PD has the scribing mark saying "A".

Attachment 2: This diode had no glass cap on it. The photodiode sensitive element is still intact. For ease of handling, it should be kept in a cage. There are four cages in the OMC lab, but they are ocuppied with the High QE PDs and others. So, the cage for this PD was offered by Rich from his office, meaning the cage was not clean.

Attachment 3: The sensor side is capped by a plate. This cap can be removed by unscrewing the two cap screws in the photo.

Attachment 4: The PD legs are shorted. (Just to match the style with the LLO one).

Attachment 5: Wrapped with AL foil and double bagged. (Repeat: It is not anything clean.)

Attachment 6: The bag was left on Rich's desk.

Attachment 1: IMG_2826.JPG  1.870 MB  | Show | Hide all | Show all
Attachment 2: IMG_2835.JPG  4.241 MB  | Hide | Hide all | Show all
IMG_2835.JPG
Attachment 3: IMG_2833.JPG  1.636 MB  | Show | Hide all | Show all
Attachment 4: IMG_2836.JPG  1.488 MB  | Hide | Hide all | Show all
IMG_2836.JPG
Attachment 5: IMG_2837.JPG  1.584 MB  | Show | Hide all | Show all
Attachment 6: DSC_0546.JPG  279 kB  | Show | Hide all | Show all
ELOG V3.1.3-